# Thermodynamic speed limit for non-adiabatic work and its classical-quantum decomposition

### v1

@@ I Introduction
Quantifying the universal bound on mechanical work is indispensable for optimization of efficient energy conversion, extraction, and storage in various thermodynamic devices, including molecular machines [1, 2], heat engines [1, 3, 4, 5], information engines [6], and quantum batteries [7]. According to the second law of thermodynamics, the work cost is minimized by adiabatically slow processes in which the system is always at equilibrium. However, realistic systems typically operate beyond the adiabatic regime and are out of equilibrium; therefore, recent studies aim to develop finite-time and nonequilibrium thermodynamic relations [8, 9, 10] and optimal protocols [11, 12, 13, 14]. When the system operates quasi-adiabatically, one could utilize linear-response relations and thermodynamic metrics to explore geometric bounds and optimization methods for work in both classical and quantum systems [15, 16, 17, 18, 19].
Far beyond the adiabatic regime, there exist several frameworks, including quantum and thermodynamic speed limits [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30], thermodynamic uncertainty relations [31, 32, 33, 9], finite-time Landauer principles [34, 35, 36, 37] and thermodynamic trade-off relations [38, 39, 40, 41], which give universal bounds on the speed of probability currents and current-like observables. For such current-like thermodynamic quantities, it is also known that information-geometric quantities such as Fisher information [42] provide important thermodynamic trade-off relations [43, 44, 45, 46]. These relations clarify that larger entropy production is required to increase the speed of physical processes when the system is subject to thermal environments. However, existing speed limits and trade-off relations cannot be directly applied to work, since work is not a current-like time-independent observable. While some studies bound the amount of work based on internal energy and heat [47], or by utilizing optimal transport theory [48] for classical Langevin systems [49, 50], quantification of thermodynamic costs that constrain the speed of non-adiabatic work extraction or consumption processes, particularly in the quantum regime, has not been well established.
@@ II Setup
We assume that the quantum system of interest is weakly coupled to a heat bath at an inverse temperature $\beta$. By assuming the standard Born-Markov-Secular approximations, we arrive at the Markovian master equation of the Gorini-Kossakowski-Sudarshan-Lindblad form [51, 52, 53], given by
$\partial_{t}\rho_{t}=\mathcal{L}_{t}[\rho_{t}]=\mathcal{H}_{t}[\rho_{t}]+\mathcal{D}_{t}[\rho_{t}],$ (1)
where
@@
$\dot{W}:=\mathrm{Tr}[\rho_{t}\dot{H}_{t}],$ (6)
where $\dot{X}_{t}:=\partial_{t}X_{t}$ is a short-hand notation of the time-derivative of the quantity $X_{t}$. In the adiabatically slow driving limit, the density operator of the system can be approximated by its instantaneous thermal state $\rho(t)=\rho_{t}^{\mathrm{eq}}+O(\dot{\lambda}_{t})$. In this regime, the work reduces to the adiabatic work, defined as
@@
$\dot{W}=\dot{W}_{\rm ad}+\mathrm{Tr}[\dot{H}_{t}\mathcal{L}_{t}^{+}[\partial_{t}\rho_{t}^{\mathrm{eq}}]]+O(\dot{\lambda}_{t}^{3}),$ (8)
where $\mathcal{L}_{t}^{+}$ is the Drazin inverse of the Loiuville superoperator $\mathcal{L}_{t}$, defined as [54, 16, 55]
@@
If $\mathcal{L}_{t}$ is diagonalizable, let us denote it as $\mathcal{L}_{t}=\sum_{\Gamma}\Gamma P_{\Gamma}$, where $\Gamma$ and $P_{\Gamma}$ are the eigenvalues and projection superoperators, respectively. Then, the Drazin inverse $\mathcal{L}_{t}^{+}$ can be expressed as $\mathcal{L}_{t}^{+}=\sum_{\Gamma\neq 0}\Gamma^{-1}P_{\Gamma}$, where $\Gamma^{-1}$ quantifies effective relaxation time-scales for each eigenmode.
It should be noted that the second term on the right-hand side of Eq. (8) can be expressed as a thermodynamic metric in the control parameter space $\lambda_{t}$, and it is possible to optimize the work by finding the optimal driving protocol $\lambda_{t}$ [16]. However, when driving is beyond this near-adiabatic regime, we need a general framework that quantifies the non-adiabatic contribution to the work. To this end, we employ techniques developed in the study of thermodynamic speed limits and derive a general upper bound on the non-adiabatic work $\dot{W}-\dot{W}_{\rm ad}$.
@@ III Main result 1: thermodynamic speed limits for non-adiabatic work
$\rho_{t}-\rho_{t}^{\mathrm{eq}}=\mathcal{L}_{t}^{+}[\mathcal{L}_{t}[\rho_{t}]]=\mathcal{L}_{t}^{+}[\partial_{t}\rho_{t}],$ (10)
@@
$\dot{W}-\dot{W}_{\rm ad}=\mathrm{Tr}[\dot{H}_{t}\mathcal{L}_{t}^{+}[\partial_{t}\rho_{t}]]=\mathrm{Tr}[\tilde{\mathcal{L}}_{t}^{+}[\dot{H}_{t}]\partial_{t}\rho_{t}].$ (11)
@@
$|\dot{W}-\dot{W}_{\rm ad}|\leq\sqrt{4g_{tt}^{\rm QF}V_{\rho_{t}}(\tilde{\mathcal{L}}_{t}^{+}[\dot{H}_{t}])}+\sqrt{2A(\rho_{t})\dot{\sigma}}=:\mathcal{B}_{1},$ (14)
which is the first main result of this paper. The proof of Eq. (14) is shown in Appendix B, where we utilize the techniques developed in Refs. [24, 30]. In the following, we explain each term that appears in Eq. (14).
  • The quantity $g^{\rm QF}_{tt}$ describes a purely quantum contribution to the quantum Fisher information metric and is defined as [56, 22]
@@
  • The quantity $V_{\rho_{t}}(\tilde{\mathcal{L}}_{t}^{+}[\dot{H}_{t}])$ quantifies the fluctuation of ~$\tilde{\mathcal{L}}_{t}^{+}[\dot{H}_{t}]$, where the variance is defined as $V_{\rho_{t}}[X]:=\mathrm{Tr}[X^{2}\rho_{t}]-(\mathrm{Tr}[X\rho_{t}])^{2}$. When $\mathcal{L}_{t}$ is diagonalizable, ~$\tilde{\mathcal{L}}_{t}^{+}$ rescales $\dot{H}_{t}$ by dividing it with effective relaxation rates $\Gamma$ for each eigenmode, meaning that ~$\tilde{\mathcal{L}}_{t}^{+}[\dot{H}_{t}]$ characterizes the relative time-scale between the driving speed $\dot{\lambda}_{t}$ of the Hamiltonian and the effective relaxation rates $\Gamma$. The analytical expression of ~$\tilde{\mathcal{L}}_{t}^{+}[\dot{H}_{t}]$ for a two-level system example is given in Eq. (92).
@@
    quantifies the number of total jumps per unit time and is called (dynamical) activity [23, 24]. Therefore, the term $A$ is an activity-like quantity that quantifies the instantaneous fluctuation of ~$\tilde{\mathcal{L}}_{t}^{+}[\dot{H}_{t}]$ multiplied by the time-scale of quantum jumps for each eigenbasis $|n(t)\rangle$ of $\rho_{t}$.
@@
    where $S(\rho):=-\mathrm{Tr}[\rho\ln\rho]$ is the von Neumann entropy of the system, $\dot{Q}:=\mathrm{Tr}[H_{t}\partial_{t}\rho_{t}]$ is the heat that quantifies the energy exchange between the system and the heat bath, and the nonnegativity of $\dot{\sigma}$ represents the second law of thermodynamics [57, 58]. The entropy production quantifies the thermodynamic cost of non-adiabatic processes and vanishes in the adiabatic limit.
The obtained thermodynamic speed limit (14) is valid for any driving speed and represents a general upper bound on non-adiabatic work. The bound depends on the fluctuation of ~$\tilde{\mathcal{L}}_{t}^{+}[\dot{H}_{t}]$ through the quantities $V_{\rho_{t}}(\tilde{\mathcal{L}}_{t}^{+}[\dot{H}_{t}])$ and $A(\rho_{t})$, where ~$\tilde{\mathcal{L}}_{t}^{+}[\dot{H}_{t}]$ characterizes the ratio between the driving speed and the effective relaxation rates. In the adiabatic limit, both $g_{tt}^{\rm QF}$ and $\dot{\sigma}$ vanish, thereby quantifying non-adiabatic costs for generating coherent unitary dynamics and stochastic quantum jump dynamics, respectively. Quite importantly, our setup assumes a thermodynamically consistent master equation (1), and thus thermodynamic costs such as the entropy production and activity-like quantity set the time-scale of the system [23, 24]. The first term on the right-hand side of (14) depends on $g_{tt}^{\rm QF}$ and therefore vanishes in the limit where Eq. (1) can be approximated by a classical master equation. On the other hand, the second term on the right-hand side of (14) does not vanish in the classical limit and represents the thermodynamic speed limit for non-adiabatic work in classical stochastic systems. Therefore, the first and second terms of the bound (14) can be interpreted as the quantum and classical part of the speed limit for non-adiabatic work, respectively. In the next section, we generalize this classical-quantum interpretation by deriving another bound based on the classical-quantum decomposition of the non-adiabatic work.
@@
In this section, we decompose the non-adiabatic work into classical and quantum contributions and derive thermodynamic speed limits for individual terms.
@@ IV.1 Classical-quantum decomposition of the non-adiabatic work
The non-adiabatic work can be decomposed into classical and quantum contributions such as $\dot{W}=\dot{W}_{\rm cl}+\dot{W}_{\rm qm}$, where
$\dot{W}_{\rm cl}:=\sum_{n}\dot{\epsilon}_{n}\langle\epsilon_{n}|\rho_{t}|\epsilon_{n}\rangle=\mathrm{Tr}[\dot{H}_{t}^{\rm cl}\rho_{t}^{\rm diag}]$ (20)
is the work originating from the time variation of the energy eigenvalues $\epsilon_{n}(t)$, and depends only on the diagonal component of the density operator, i.e., $\rho_{t}^{\rm diag}=\sum_{n}|\epsilon_{n}\rangle\langle\epsilon_{n}|\rho_{t}|\epsilon_{n}\rangle\langle\epsilon_{n}|$. Therefore, Eq. (20) quantifies the classical contribution to the work. On the other hand,
$\dot{W}_{\rm qm}:=\sum_{n}\epsilon_{n}\Bigl{(}\langle\dot{\epsilon}_{n}|\rho_{t}|\epsilon_{n}\rangle+\langle\epsilon_{n}|\rho_{t}|\dot{\epsilon}_{n}\rangle\Bigr{)}=\mathrm{Tr}[\dot{H}_{t}^{\rm qm}\delta\rho_{t}^{\rm coh}]$ (21)
depends solely on the coherence of the density operator $\delta\rho_{t}^{\rm coh}:=\rho_{t}-\rho_{t}^{\rm diag}$ and thus quantifies quantum contribution to the work. Here, we have also decomposed the time derivative of the Hamiltonian into the classical and quantum contributions, defined by
@@
We note that in the adiabatic limit, we have $\dot{W}_{\rm cl}=\dot{W}_{\rm ad}+O(\dot{\lambda}_{t}^{2})$ and $\dot{W}_{\rm qm}=O(\dot{\lambda}_{t}^{2})$. Therefore, the non-adiabatic contribution to the work is decomposed as

### v2

@@ I Introduction
Quantifying the universal bound on mechanical work is indispensable for optimization of efficient energy conversion, extraction, and storage in various thermodynamic devices, including molecular machines Seifert (2012); Leighton and Sivak (2024), heat engines Seifert (2012); del Campo et al. (2018); Myers et al. (2022); Cangemi et al. (2024), information engines Parrondo et al. (2015), and quantum batteries Campaioli et al. (2024). According to the second law of thermodynamics, the work cost is minimized by adiabatically slow processes in which the system is always at equilibrium. However, realistic systems typically operate beyond the adiabatic regime and are out of equilibrium; therefore, recent studies aim to develop finite-time and nonequilibrium thermodynamic relations Seifert (2019); Horowitz and Gingrich (2020); Wolpert et al. (2024) and optimal protocols Kosloff and Rezek (2017); Deffner and Bonança (2020); Guéry-Odelin et al. (2023); Blaber and Sivak (2023). When the system operates quasi-adiabatically, one could utilize linear-response relations and thermodynamic metrics to explore geometric bounds and optimization methods for work in both classical and quantum systems Brandner and Saito (2020); Scandi and Perarnau-Llobet (2019); Abiuso et al. (2020); Miller et al. (2019); Rolandi et al. (2023).
Far beyond the adiabatic regime, there exist several frameworks, including quantum and thermodynamic speed limits Deffner and Campbell (2017); Gong and Hamazaki (2022); Pires et al. (2016); Shiraishi et al. (2018); Funo et al. (2019); Nicholson et al. (2020); García-Pintos et al. (2022); Ito and Dechant (2020a); Van Vu and Hasegawa (2021); Van Vu and Saito (2023); Sekiguchi et al. (2024), thermodynamic uncertainty relations Barato and Seifert (2015); Gingrich et al. (2016a); Horowitz and Gingrich (2017, 2020), finite-time Landauer principles Goold et al. (2015); Proesmans et al. (2020a); Zhen et al. (2021); Van Vu and Saito (2022) and thermodynamic trade-off relations Gingrich et al. (2016b); Shiraishi et al. (2016); Tajima and Funo (2021); Yoshimura et al. (2025), which give universal bounds on the speed of probability currents and current-like observables. For such current-like thermodynamic quantities, it is also known that information-geometric quantities such as Fisher information Cover (1999) provide important thermodynamic trade-off relations Dechant (2018); Hasegawa and Van Vu (2019a); Ito and Dechant (2020b); Otsubo et al. (2020). These relations clarify that larger entropy production is required to increase the speed of physical processes when the system is subject to thermal environments. However, existing speed limits and trade-off relations cannot be directly applied to work, since work is not a current-like time-independent observable. While some studies bound the amount of work based on internal energy and heat Aghion and Green (2023), or by utilizing optimal transport theory Benamou and Brenier (2000) for classical Langevin systems Aurell et al. (2011); Proesmans et al. (2020b), quantification of thermodynamic costs that constrain the speed of non-adiabatic work extraction or consumption processes, particularly in the quantum regime, has not been well established.
@@ II Setup
We assume that the quantum system of interest is weakly coupled to a heat bath at an inverse temperature $\beta$. By assuming the standard Born-Markov-Secular approximations, we arrive at the Markovian master equation of the Gorini-Kossakowski-Sudarshan-Lindblad form V. Gorini and Sudarshan (2008); Lindblad (1976); Breuer and Petruccione (2002), given by
$\dot{\rho}_{t}=\mathcal{L}_{t}[\rho_{t}]=\mathcal{H}_{t}[\rho_{t}]+\mathcal{D}_{t}[\rho_{t}],$ (1)
where $\dot{X}_{t}:=\partial_{t}X_{t}$ is a short-hand notation of the time-derivative of the quantity $X_{t}$, and
@@
$\dot{W}:=\mathrm{Tr}[\rho_{t}\dot{H}_{t}].$ (6)
In the adiabatically slow driving limit, the density operator of the system can be approximated by its instantaneous thermal state $\rho(t)=\rho_{t}^{\mathrm{eq}}+O(\dot{\lambda}_{t})$. In this regime, the work reduces to the adiabatic work, defined as
@@
$\dot{W}=\dot{W}_{\rm ad}+\mathrm{Tr}[\dot{H}_{t}\mathcal{L}_{t}^{+}[\dot{\rho}_{t}^{\mathrm{eq}}]]+O(\dot{\lambda}_{t}^{3}),$ (8)
where $\mathcal{L}_{t}^{+}$ is the Drazin inverse of the Liouville superoperator $\mathcal{L}_{t}$, defined as Drazin (1958); Scandi and Perarnau-Llobet (2019); Chen et al. (2021)
@@
If $\mathcal{L}_{t}$ is diagonalizable, let us denote it as $\mathcal{L}_{t}=\sum_{\Gamma}\Gamma P_{\Gamma}$, where $\Gamma$ and $P_{\Gamma}$ are the eigenvalues and projection superoperators, respectively. Then, the Drazin inverse $\mathcal{L}_{t}^{+}$ can be expressed as $\mathcal{L}_{t}^{+}=\sum_{\Gamma\neq 0}\Gamma^{-1}P_{\Gamma}$, where the real part of $-\Gamma^{-1}$ quantifies the relaxation time-scales for each eigenmode.
It should be noted that the second term on the right-hand side of Eq. (8) can be expressed as a thermodynamic metric in the control parameter space $\lambda_{t}$, and it is possible to optimize the work by finding the optimal driving protocol $\lambda_{t}$ Scandi and Perarnau-Llobet (2019). However, when driving is beyond this near-adiabatic regime, we need a general framework that quantifies the non-adiabatic contribution to the work. To this end, we employ techniques developed in the study of thermodynamic speed limits and derive a general upper bound on the non-adiabatic work $\dot{W}-\dot{W}_{\rm ad}$.
@@ III Main result 1: thermodynamic speed limits for non-adiabatic work
$\rho_{t}-\rho_{t}^{\mathrm{eq}}=\mathcal{L}_{t}^{+}[\mathcal{L}_{t}[\rho_{t}]]=\mathcal{L}_{t}^{+}[\dot{\rho}_{t}],$ (10)
@@
$\dot{W}-\dot{W}_{\rm ad}=\mathrm{Tr}[\dot{H}_{t}\mathcal{L}_{t}^{+}[\dot{\rho}_{t}]]=\mathrm{Tr}[\tilde{\mathcal{L}}_{t}^{+}[\dot{H}_{t}]\dot{\rho}_{t}].$ (11)
@@
$|\dot{W}-\dot{W}_{\rm ad}|\leq\underbrace{\sqrt{4g_{tt}^{\rm QF}V_{\rho_{t}}(\tilde{\mathcal{L}}_{t}^{+}[\dot{H}_{t}])}}_{\text{non-diagonal}}+\underbrace{\sqrt{2A(\rho_{t})\dot{\sigma}}}_{\text{diagonal}}=:\mathcal{B}_{1}^{\rm nd}+\mathcal{B}_{1}^{\rm d},$ (14)
which is the first main result of this paper. The proof of Eq. (14) is shown in Appendix B, where we utilize the techniques developed in Refs. Funo et al. (2019); Sekiguchi et al. (2024). Here, the bound contains two terms $\mathcal{B}_{1}^{\rm nd}$ and $\mathcal{B}_{1}^{\rm d}$, depending on the non-diagonal and diagonal contributions using the basis of the density operator $\rho_{t}$, respectively. In the following, we explain each term that appears in Eq. (14).
  • The quantity $g^{\rm QF}_{tt}$ describes a purely quantum contribution to the quantum Fisher information metric and is defined as Braunstein and Caves (1994); Pires et al. (2016)
@@
  • The quantity $V_{\rho_{t}}(\tilde{\mathcal{L}}_{t}^{+}[\dot{H}_{t}])$ quantifies the fluctuation of ~$\tilde{\mathcal{L}}_{t}^{+}[\dot{H}_{t}]$, where the variance is defined as $V_{\rho_{t}}[X]:=\mathrm{Tr}[X^{2}\rho_{t}]-(\mathrm{Tr}[X\rho_{t}])^{2}$. When $\mathcal{L}_{t}$ is diagonalizable, ~$\tilde{\mathcal{L}}_{t}^{+}$ rescales $\dot{H}_{t}$ by dividing it with effective relaxation rates $\Gamma$ for each eigenmode, meaning that ~$\tilde{\mathcal{L}}_{t}^{+}[\dot{H}_{t}]$ characterizes the relative time-scale between the driving speed $\dot{\lambda}_{t}$ of the Hamiltonian and the relaxation rates $\Gamma$. The analytical expression of ~$\tilde{\mathcal{L}}_{t}^{+}[\dot{H}_{t}]$ for a two-level system example is given in Eq. (92).
@@
    quantifies the number of total jumps per unit time and is called (dynamical) activity Shiraishi et al. (2018); Funo et al. (2019). Therefore, the term $A$ is an activity-like quantity that quantifies the instantaneous fluctuation of ~$\tilde{\mathcal{L}}_{t}^{+}[\dot{H}_{t}]$ multiplied by the time-scale of quantum jumps for each eigenbasis $|n(t)\rangle$ of $\rho_{t}$. See also Eq. (42) for the simplified expression of $A$ in the case of two-level systems.
@@
    where $S(\rho):=-\mathrm{Tr}[\rho\ln\rho]$ is the von Neumann entropy of the system, $\dot{Q}:=\mathrm{Tr}[H_{t}\dot{\rho}_{t}]$ is the heat that quantifies the energy exchange between the system and the heat bath, and the nonnegativity of $\dot{\sigma}$ represents the second law of thermodynamics Spohn (1978); Funo et al. (2018). The entropy production quantifies the thermodynamic cost of non-adiabatic processes and vanishes in the adiabatic limit.
The obtained thermodynamic speed limit (14) is valid for any driving speed and represents a general upper bound on non-adiabatic work. In the adiabatic limit, both $g_{tt}^{\rm QF}$ and $\dot{\sigma}$ vanish, thereby the bound quantifies geometric and thermodynamic costs to produce work in the non-adiabatic regime.
The non-diagonal part of the bound $\mathcal{B}_{1}^{\rm nd}$ depends on $g_{tt}^{\rm QF}$, which characterizes the speed of coherent unitary part of the time-evolution. Moreover, it vanishes in the limit where Eq. (1) can be approximated by a classical master equation. On the other hand, $\mathcal{B}_{1}^{\rm d}$ depends on the entropy production $\dot{\sigma}$ and the activity-like quantity $A$, which quantifies the speed of the stochastic jump part of the time-evolution Shiraishi et al. (2018); Funo et al. (2019). Moreover, it vanishes in the limit where Eq. (1) can be approximated by the time-evolution for isolated quantum systems. Based on the above discussion, the bounds $\mathcal{B}_{1}^{\rm nd}$ and $\mathcal{B}_{1}^{\rm d}$ can be interpreted as the quantum and classical part of the speed limit for non-adiabatic work, respectively. In the next section, we generalize this classical-quantum interpretation by deriving another bound based on the classical-quantum decomposition of the non-adiabatic work.
@@
In this section, we decompose the non-adiabatic work into classical and quantum contributions and derive thermodynamic speed limits for individual terms. Specifically, instead of splitting the bound into diagonal and non-diagonal contributions using the basis of $\rho_{t}$ as done in Sec. III, we now use the basis of $H_{t}$ and derive a similar bound.
@@ IV.1 Classical-quantum decomposition of the non-adiabatic work
To begin with, we decompose the time derivative of the Hamiltonian into the classical and quantum contributions, defined by
@@
Based on the decomposition (20), the non-adiabatic work can be similarly decomposed into classical and quantum contributions such as $\dot{W}=\dot{W}_{\rm cl}+\dot{W}_{\rm qm}$, where
$\dot{W}_{\rm cl}:=\sum_{n}\dot{\epsilon}_{n}\langle\epsilon_{n}|\rho_{t}|\epsilon_{n}\rangle=\mathrm{Tr}[\dot{H}_{t}^{\rm cl}\rho_{t}^{\rm cl}]$ (21)
is the work originating from the time variation of the energy eigenvalues $\epsilon_{n}(t)$, and depends only on the diagonal component of the density operator using the energy eigenbasis, i.e., $\rho_{t}^{\rm cl}=\sum_{n}|\epsilon_{n}\rangle\langle\epsilon_{n}|\rho_{t}|\epsilon_{n}\rangle\langle\epsilon_{n}|$. Therefore, Eq. (21) quantifies the classical contribution to the work. On the other hand,
$\dot{W}_{\rm qm}:=\sum_{n}\epsilon_{n}\Bigl{(}\langle\dot{\epsilon}_{n}|\rho_{t}|\epsilon_{n}\rangle+\langle\epsilon_{n}|\rho_{t}|\dot{\epsilon}_{n}\rangle\Bigr{)}=\mathrm{Tr}[\dot{H}_{t}^{\rm qm}\rho_{t}^{\rm qm}]$ (22)
depends solely on the coherence of the density operator $\rho_{t}^{\rm qm}:=\rho_{t}-\rho_{t}^{\rm cl}$ and thus quantifies quantum contribution to the work. We note that in the adiabatic limit, we have $\dot{W}_{\rm cl}=\dot{W}_{\rm ad}+O(\dot{\lambda}_{t}^{2})$ and $\dot{W}_{\rm qm}=O(\dot{\lambda}_{t}^{2})$. Therefore, the non-adiabatic contribution to the work is decomposed as